OpenStudy (anonymous):

A sphere has a center (2,3,5) and has a radius of 3. Find the equation of all planes containing the points (4,0,0) and (0,0,4) that are tangent to this circle. Hint: Let (a,b,c) be the point on the plane that touches the sphere. FML

OpenStudy (anonymous):

This is a tough problem, but I can give some hints

OpenStudy (anonymous):

A point where it's tangent is a solution to both the equation of the plane and the sphere, AND both a solution to the equation to the plane and the sphere's partial derivatives.

OpenStudy (anonymous):

Alternatively, you can use the normal vectors of the plane and sphere. The normal of a sphere is simply the point itself (scaled possibly), and the plane is going to be simple to calculate. You just have to do it arbitrarily.

OpenStudy (anonymous):

Right now all I have is the equation of the sphere and the normal vector which is <2-a, 3-b, 5-c>. I can not find (a,b,c) I don't know how I can go about solving for each letter to get that point.

OpenStudy (anonymous):

I tried dot product or just making the magnitude of the normal vector equal to the radius of the sphere but it becomes a three unknown system of equations and I can't solve that.

OpenStudy (anonymous):

So, try this. Let (a,b,c) be the point on the plane that touches the sphere. Then the normal vector of the plane is the line from the center of the sphere to (a,b,c), although you might have to scale this to get absolute value of 1 for your vector. Then you have the normal vector for the plane. You also have some points on the plane. You should be able to find an equation.

OpenStudy (anonymous):

I think that's how you do it actually.

OpenStudy (anonymous):

That's what I did but it still leaves me with three unknowns. My prof said at the end the equation of the planes should not have any a, b, or c's.

OpenStudy (anonymous):

The equation of my plane would be (2-a)x + (3-b) + (5-c)z = 4 - 4a using the point (4,0,0)

OpenStudy (anonymous):

That's odd, since it asked for the equation of all planes. How can you have no free variables?

OpenStudy (anonymous):

Well you can just not a, b, and c. X, Y, and Z are fine because they are part of the equation of the plane. basically he wants us the fund the point (a,b,c) and I don't know how to find it with the given information.

OpenStudy (anonymous):

Ah, here's how: There are only finitely many planes?

OpenStudy (anonymous):

Yes. Well at least I believe so.

OpenStudy (anonymous):

Because my visualization is telling me there are only two solutions.

OpenStudy (anonymous):

I thought so two, to be completely honest I don't know.

OpenStudy (anonymous):

Btw thank you so much for your help, I really appreciate all your input.

OpenStudy (anonymous):

Ya. This is the most interesting problem I've seen tonight.

OpenStudy (anonymous):

Yeah, I don't know why my prof does this to us -.- The only thing he told me was that the normal vector was correct, <2-a, 3-b, 5-c>, and then he told me to get a second vector from the plane which I got from point (4,0,0) to the point (a,b,c) which is <4-a,-b,-c>.

OpenStudy (anonymous):

Ah, so this was where you tried solving for the dot product = 0?

OpenStudy (anonymous):

Exactly, they are perpendicular so there for the dot product equals zero but it still leaves me with three unknowns.

OpenStudy (anonymous):

I see that ;(

OpenStudy (anonymous):

I guess I will just wait until tomorrow, to go to a tutoring session before it's due.

OpenStudy (anonymous):

Good luck

OpenStudy (anonymous):

Thanks, and thank you for all your input!

OpenStudy (anonymous):

Did you try using the other vector from the plane as well?

OpenStudy (anonymous):

<2-a, 3-b, 5-c> <4-a,-b,-c> = 0 = <2-a, 3-b, 5-c> <-a,-b,4-c>. That should kill one of the unknowns I'd hope.

OpenStudy (anonymous):

anyway, good luck with this.

OpenStudy (anonymous):

So, my prof gave us another hint he said that you need 3 equations to solve for all three unknowns. So I have the equation of the sphere at point (a,b,c) which is (a-2)^2 + (b-3)^2 + (c-5)^2 = 9 Also have the dot product which equals zero which is: a^2 - 6a + b^2 -3b +c^2 - 5c = 0 Now I just need one more equation, also he told us that both points (4,0,0) and (0,0,4) are a part of both planes that I am solving for.

OpenStudy (anonymous):

Okay so I found the other equation for the system of equations which is: a^2 -2a + b^2 -3b + c^2 -9c = -20 and I found c = 3+a

OpenStudy (phi):

I am curious to see how you get the answer. I found one generic way which involved using a rotation matrix, and a slightly simpler way for this specific problem. I did not see a nice way to solve for (a,b,c) algebraically... when you find the answer, please post it.